OpenStudy (anonymous):

Find the x-intercept of the line whose equation is 8x + 2y = 4.

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

just set y=0 8x+2(0)=4 8x=4 x=4/8 x=1/2